# Compact kinks in a modified Christ-Lee model

### v1

@@ VI Conclusions
For all mentioned cases, static solutions for compact kinks have been obtained numerically. We have found that for the cases of small $\epsilon$, the kink in the $(-1,1)$ sector can be interpreted as composed of two subkinks in the $(-1,0)$ and $(0,1)$ sectors. We have also analyzed the dynamics of such kinks under small perturbations by finding its internal modes. As usual in the case of compact kinks, the stability potential for small oscillatory perturbations features infinite barriers at the borders of the compact kinks, resulting in the internal modes to also be compact. For $\epsilon=0$, we have found that the presence of the subkinks makes the modes’ frequency spectrum degenerate. However, the degeneracy is quickly broken when increasing values of $\epsilon$.
The interactions between the compact kinks have also been studied through a careful numerical treatment of the compactons collision. We have found that the parameter $n$ is related to the emission of radiation in kink-kink collisions in the $(-1,0)+(0,1)$ sector, with larger amounts of radiation emitted when $n$ is large, while the $n=1/2$ case has no radiation. In the kink-antikink collisions in the $(0,1)+(1,0)$ and $(1,0)+(0,1)$ sectors, the parameter $n$ has greatly influenced the critical velocity separating the capture and escape cases, as well as the formation of large central oscillons and the size of the oscillons that compose the radiation.
In the case of $(-1,1)+(1,-1)$ collisions, we have showed that, for small values of $\epsilon$, the internal structure of the kink as two subkinks plays an important role. Indeed, for $\epsilon$ small, we have detected large windows of velocities for which the subkinks decouple from the larger kink, leading to the formation of false vacuum bubbles. Even the cases without decoupling, the more rich inner structure of the $(-1,1)$ kink renders the collision dynamics more complicated, causing the appearance of extra capture windows in the $n=1/2$, $\epsilon=0$ case, as well as a window of multi-bounce followed by escape in the $n=1/2$, $\epsilon=0.1$ case.
Perhaps somewhat surprisingly, also in the simpler double well case ($\epsilon\rightarrow\infty$), we have found that the kink-antikink collision dynamics is less straightforward than previously thought. The $n$ parameter has been shown to greatly influence the dynamics. For example, the $n=1/2$ case displays almost no escape cases, except for a small velocities window, for which escape occurs only after a few bounces. Cases of multi-bounce followed by escape mostly disappear when we increase $n$ to $1$, but reappear for larger values of $n$.
@@ Acknowledgements.
This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior – Brasil (CAPES) – Finance Code 001. The authors would like to thank the Centro de Computação Avançada e Multidisciplinar from Universidade Estadual de Santa Cruz for their cooperation, provision and operation of the computing facilities, and their commitment to investing in research, development and innovation for education, industry and society as a whole.
@@ Appendix A Numerical methods
For the scans of multiple simulations with $\epsilon=0$ and $\epsilon=\infty$, such as the ones in figures 21 and 26, we have adopted a different strategy. Since the scans only use the field at $x=0$, we have performed the simulations only for $x\in[-x_{\text{max}},x_{\text{max}}]$, with $x_{\text{max}}$ chosen as $12$ for the $(1,0)+(0,1)$ and $(0,1)+(1,0)$ cases, and as $15$ for the $(-1,1)+(1,-1)$ case with $\epsilon\in\{0,\infty\}$. To guarantee that the borders would not play a role in the simulation outcome, we have introduced a dissipative term to the field equation at a region of size $l_{D}$ before each border. That means that, for $x\in[x_{\text{max}}-l_{D},x_{\text{max}}]$ and $x\in[-x_{\text{max}},-x_{\text{max}}+l_{D}]$, the actual simulation was performed with the field equation
@@
We have chosen $l_{D}=2$ and $\alpha_{D}=2$ for the $(1,0)+(0,1)$ and $(0,1)+(1,0)$ cases, and $l_{D}=5$ and $\alpha_{D}=0.5$ for the $(-1,1)+(1,-1)$ case, with $\epsilon\in\{0,\infty\}$. For the $(-1,1)+(1,-1)$ case with $0<\epsilon<\infty$, we have performed the simulation in the full interval without using the dissipative term due to the formation of large false vacuum bubbles.
The introduction of the dissipative term near the borders guarantees that outgoing structures will not be reflected back to the region of interest. The smaller grid also meant that we had fewer degrees of freedom that could possibly accumulate numerical error, allowing us to use step sizes as large as $\Delta x=0.02$. The introduction of the dissipative borders and the increased step sizes did not change the scan results while severely reducing the required computation time. We further improved the computational performance by exploring the symmetry of the initial conditions and only simulating the region $x\geq 0$. This numerical setup is similar to the one used in ref. [37].

### v2

@@ VI Conclusions
For all mentioned cases, static solutions for compact kinks have been obtained numerically. We have found that for the cases of small $\epsilon$, the kink in the (1,1)(-1,1)( - 1 , 1 ) sector can be interpreted as composed of two subkinks in the (1,0)(-1,0)( - 1 , 0 ) and (0,1)(0,1)( 0 , 1 ) sectors. We have also analyzed the dynamics of such kinks under small perturbations by finding its internal modes. As usual in the case of compact kinks, the stability potential for small oscillatory perturbations features infinite barriers at the borders of the compact kinks, resulting in the internal modes to also be compact. For $\epsilon=0$, we have found that the presence of the subkinks makes the modes’ frequency spectrum degenerate. However, the degeneracy is quickly broken when increasing values of $\epsilon$.
The interactions between the compact kinks have also been studied through a careful numerical treatment of the compactons collision. We have found that the parameter $n$ is related to the emission of radiation in kink-kink collisions in the (1,0)+(0,1)(-1,0)+(0,1)( - 1 , 0 ) + ( 0 , 1 ) sector, with larger amounts of radiation emitted when $n$ is large, while the $n=1/2$ case has no radiation. In the kink-antikink collisions in the (0,1)+(1,0)(0,1)+(1,0)( 0 , 1 ) + ( 1 , 0 ) and (1,0)+(0,1)(1,0)+(0,1)( 1 , 0 ) + ( 0 , 1 ) sectors, the parameter $n$ has greatly influenced the critical velocity separating the capture and escape cases, as well as the formation of large central oscillons and the size of the oscillons that compose the radiation.
In the case of $+(1,-1)$ ( 1 , - 1 ) collisions, we have showed that, for small values of $\epsilon$, the internal structure of the kink as two subkinks plays an important role. Indeed, for $\epsilon$ small, we have detected large windows of velocities for which the subkinks decouple from the larger kink, leading to the formation of false vacuum bubbles. Even the cases without decoupling, the more rich inner structure of the (1,1)(-1,1)( - 1 , 1 ) kink renders the collision dynamics more complicated, causing the appearance of extra capture windows in the $n=1/2$, $\epsilon=0$ case, as well as a window of multi-bounce followed by escape in the $n=1/2$, $\epsilon=0.1$ case.
Perhaps somewhat surprisingly, also in the simpler double well case ($\epsilon\rightarrow\infty$), we have found that the kink-antikink collision dynamics is less straightforward than previously thought. The $n$ parameter has been shown to greatly influence the dynamics. For example, the $n=1/2$ case displays almost no escape cases, except for a small velocities window, for which escape occurs only after a few bounces. Cases of multi-bounce followed by escape mostly disappear when we increase $n$ to 111, but reappear for larger values of $n$.
@@ Acknowledgements.
We thank P. Klimas and J. S. Streibel for helpful comments. This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior – Brasil (CAPES) – Finance Code 001. The authors would like to thank the Centro de Computação Avançada e Multidisciplinar from Universidade Estadual de Santa Cruz for their cooperation, provision and operation of the computing facilities, and their commitment to investing in research, development and innovation for education, industry and society as a whole.
@@ Appendix A Numerical methods
For the scans of multiple simulations with $\epsilon=0$ and $\epsilon=\infty$, such as the ones in figures 21 and 26, we have adopted a different strategy. Since the scans only use the field at $x=0$, we have performed the simulations only for $x\in[-x_{\text{max}},x_{\text{max}}]$, with $x_{\text{max}}$ chosen as 121212 for the (1,0)+(0,1)(1,0)+(0,1)( 1 , 0 ) + ( 0 , 1 ) and (0,1)+(1,0)(0,1)+(1,0)( 0 , 1 ) + ( 1 , 0 ) cases, and as 151515 for the $+(1,-1)$ ( 1 , - 1 ) case with $\epsilon\in\{0,\infty\}$. To guarantee that the borders would not play a role in the simulation outcome, we have introduced a dissipative term to the field equation at a region of size $l_{D}$ before each border. That means that, for $x\in[x_{\text{max}}-l_{D},x_{\text{max}}]$ and $x\in[-x_{\text{max}},-x_{\text{max}}+l_{D}]$, the actual simulation was performed with the field equation
@@
We have chosen $l_{D}=2$ and $\alpha_{D}=2$ for the (1,0)+(0,1)(1,0)+(0,1)( 1 , 0 ) + ( 0 , 1 ) and (0,1)+(1,0)(0,1)+(1,0)( 0 , 1 ) + ( 1 , 0 ) cases, and $l_{D}=5$ and $\alpha_{D}=0.5$ for the $+(1,-1)$ ( 1 , - 1 ) case, with $\epsilon\in\{0,\infty\}$. For the $+(1,-1)$ ( 1 , - 1 ) case with $0<\epsilon<\infty$, we have performed the simulation in the full interval without using the dissipative term due to the formation of large false vacuum bubbles.
The introduction of the dissipative term near the borders guarantees that outgoing structures will not be reflected back to the region of interest. The smaller grid also meant that we had fewer degrees of freedom that could possibly accumulate numerical error, allowing us to use step sizes as large as $\Delta x=0.02$. The introduction of the dissipative borders and the increased step sizes did not change the scan results while severely reducing the required computation time. The dissipative regions were found to be very effective. The radiation energy reaching it quickly vanishes, while the solitons stop moving before reaching the borders of the simulated region. We further improved the computational performance by exploring the symmetry of the initial conditions and only simulating the region $x\geq 0$.
The chosen numerical setup, actually similar to the one used in ref. [35], has shown the best performance among other ones with comparable accuracy. We controlled the accuracy by simulating standard configurations, such as uniformly moving kinks and signum-Gordon oscillons, and comparing the results with the known solutions. Despite the non-analytic nature of the potential, in all tested cases, the simulation results reproduced the known solutions within the method accuracy. Note that the non-analytic potential can be seen as the limit of infinite mass of a more usual potential [16, 18], and the compact kinks are the limiting case of kinks with an increasingly fast approach to the vacuum. Therefore, any theoretical constraint on the application of the numerical methods could be circumvented by working with a smoothed out potential with high enough mass. We also checked energy conservation for the cases without the dissipative term and the total energy of the system did not deviate from the initial amount by more than 5%.